OpenStudy (anonymous):

Danielle has a CD at Crossland Bank. She invests 21,350 for five years at 4.75% interest, compounded quarterly. What is her ending balance?

OpenStudy (anonymous):

@amistre64

OpenStudy (mathmale):

My standard response to questions like these is: "What's the formula for compound interest?" Mind typing that in? Here, A = Amount (accumulation) P= Principal (starting amount) N= number of times interest is paid/compounded per year t= number of years during which interest accumulates

OpenStudy (mathmale):

r = interest rate, as a decimal fraction

OpenStudy (anonymous):

I've tried three times with the formula.

OpenStudy (mathmale):

Using a calculator? What kind?

OpenStudy (anonymous):

TI-30XS

OpenStudy (mathmale):

If you have the formula at hand, please type it out here. Then we'll go through the calculations on your calc.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I believe for this problem it would be: A=P (1+r/n)^nt?

OpenStudy (mathmale):

that's great. thanks for typing it in. Let's start with the interest rate. your rate is 4.75%. convert that to a decimal fraction and then divide the result by 4.

OpenStudy (mathmale):

We divide by 4 because compounding occurs quarterly here.

OpenStudy (anonymous):

A=245,000(1+0.69)^12?

OpenStudy (anonymous):

Lol, now I see what I did wrong, nevermind. Thank you.

OpenStudy (mathmale):

Important: Note that we've started with the stuff INSIDE parentheses. Assume you're familiar with "order of operations" rules. Unfort., that expression is not quite correct. Pleae be patient and go thru the ops with me. r/n = 0.0475/4 = ?

OpenStudy (mathmale):

Want to try this on your own, or still interested in my help?